# Orthogonal projections of hypercubes

### v2

@@ Abstract
Projections of hypercubes are widely used for visualizing high-dimensional binary state spaces across various scientific fields. Conventional methods for projecting hypercubes, however, face practical difficulties. Manual methods require nontrivial adjustments of the projection basis, while optimization-based algorithms limit the interpretability and reproducibility of the resulting plots. These limitations motivate the exploration of theoretically analyzable projection algorithms such as principal component analysis (PCA). Here, we investigate the mathematical properties of PCA for projecting hypercubes. Our numerical and analytical results show that PCA effectively captures polarized distributions within the state space. This property enables the assessment of the asymptotic distribution of projected vertices and error bounds, thus characterizing the performance of PCA in the projected space. We demonstrate the application of PCA to visualize the hypercubic energy landscapes of Ising spin systems. By adding projected hypercubic edges, these visualizations reveal pathways of correlated spin flips. Our work contributes to a better understanding of how PCA can be applied to explore high-dimensional binary data.
@@ I Introduction
Human interest in higher dimensions, particularly the fourth dimension, has been emerging since the 19th century [1, 2, 3, 4]—before the theory of relativity by A. Einstein [5, 6], which is well known as the theory of four-dimensional spacetime [7]. Unlike the theory of relativity, 19th-century mathematicians, physicists, and philosophers focused on the spatial fourth dimension. Among them, C. H. Hinton has been influential because his works [8] indicate that the tesseract, or four-dimensional cube, can embody four-dimensional space. Since then high-dimensional space has been explored with hypercubes.
@@ II Orthogonal Projections of Hypercubes
Figure 2: A gallery of orthogonal projections of hypercubes. Colored arrows represent the contribution basis corresponding to the original dimensions. The boxes on the bottom right indicate the correspondence between the colors and the original dimensions. (a) An isometric projection of a cube. Notice that $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$ are overlapped. (b) An isometric projection of a tesseract. (c) A Hamming projection of a cube. (d) A Hamming projection of a tesseract. Notice that $\begin{bmatrix}1&0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1&0\end{bmatrix}^{\top}$ are overlapped. In Hamming projections (c) and (d), the contribution basis of each dimension has the same horizontal contribution. (e) A fractal projection of a six-dimensional hypercube. (f) A fractal projection of an eight-dimensional hypercube. In fractal projections (e) and (f), the contribution basis of the first half (left half) of the code is ten times longer than the rest. More projections of hypercubes are available in the Supplemental Material [105].
@@ II.2 Contribution basis
The interpretation of the contribution basis in Eq. (2) is the contribution from the $i$th dimension to the projected space. Because $b_{i}\in\left\{1,0\right\}$, the contribution basis corresponds to the projected unit vector of each dimension. Thus, by introducing the contribution basis, we can decompose the resulting projection into the contributions from each dimension as shown in Eq. (2). The remaining question is how to determine the contribution bases $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$.
@@ II.3 Isometric projection
One particularly regular and symmetric projection is called isometric projection, where a Petrie polygon is projected as a regular polygon located at the periphery of the projection of a polytope [9]. Petrie polygons are equatorial polygons, lying in planes crossing the center of the object and inscribed in great circles of the circumsphere of the object. To achieve isometric projection, one needs the projected unit vector of the $i$th dimension to form consecutive edges of the projected Petrie polygon, i.e., $\bm{e}_{i}$ form a Petrie polygon together with a reversed vector $-\bm{e}_{i}$. To project the Petrie polygon as a regular polygon, the contribution bases are determined as $\bm{e}_{i}=\begin{bmatrix}\cos\left(\theta_{i}\right)&\sin\left(\theta_{i}\right)\end{bmatrix}^{\top}$, where $\theta_{i}=\frac{\uppi}{N}\left(i-1\right)+\phi$, with a constant $\phi$. The resulting isometric projections and vectors $\bm{e}_{i}$ used to create them are shown in Figs. 2(a) and 2(b), where Petrie polygons are projected as a regular hexagon [Fig. 2(a)] and octagon [Fig. 2(b)]. We determine the angle $\phi=\frac{\uppi}{2N}$ for symmetry of the projections. In isometric projections, all edges of the hypercube are drawn with the same length.
@@ II.4 Hamming and fractal projections
By modifying the contribution bases $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ for projection, one can view a high-dimensional hypercube from various angles. Several methods reflecting the cross-section of the hypercube have already been suggested [9]. Here, we introduce two other orthogonal projections of hypercubes used in sciences.
@@ II.5 Exchanging the labels of vertices
Here, we show a visual understanding of this conversion through contribution bases. By reversing the direction of the contribution basis, one can swap the projected coordinates of the rest, cf. Figs. 2(a) and 2(b), or Figs. 3(a) and 3(b). The dimension of the reversed contribution basis corresponds to the dimension where there is a $1$ in $\bm{b}^{\prime}\oplus\bm{b}^{\prime\prime}$. Visualizing the projected vector provides a way to understand exchanging the labels via reversing the direction of unit vectors.
Figure 3: Exchanging the labels of vertices by reversing the direction of the contribution basis of corresponding digits. The boxes on the bottom right indicate the correspondence between the colors and the dimensions. (a) By reversing the contribution basis for the second digit, one can exchange the cubic labels of pairs of vertices, $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1\end{bmatrix}^{\top}$, and $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Compare this with Fig. 2(a). (b) Obtaining different labels for a tesseract by swapping two contribution bases. The first and fourth unit vectors are reversed, cf. Fig. 2(b).
@@ II.6 The limitation of manual orthogonal projections
Although one can create reproducible and interpretable (and even visually label-exchangeable) two-dimensional projections of hypercubes in the abovementioned way, the disadvantage is that one must manually determine the contribution basis of each dimension. For low-dimensional hypercubes, it is manageable to adjust the contribution basis oneself to make the projections easy to interpret. Still, it is nontrivial to decide on contribution bases for high-dimensional hypercubes. Thus, automatic and unsupervised projection methods are demanded, especially for projections of high-dimensional hypercubes. How can one project a hypercube with such a strategy?
@@ III.2 PCA and biplot
From now on, we call $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ biplot bases. Notice that, when the data points are binary vertices $\bm{b}\in\left\{1,0\right\}^{N}$ of a system, the biplot bases exactly match the basis of the orthogonal projection of a hypercube, while when the data points are Ising vertices $\bm{s}\in\left\{+1,-1\right\}^{N}$, the biplot bases with doubled magnitudes exactly match the contribution bases. This difference arises from the difference in the length of hypercubic edges: they are one for binary vertices and two for Ising vertices. This is why we add factor $2$ in Eq. (10). By performing PCA and plotting biplot bases, one can obtain a reproducible, interpretable, and automatic projection of a hypercube.
@@ IV Hypercubic PCA
In the previous Sec. III, we introduced PCA as a method to project hypercubes and presented the interpretation of the projection using biplot bases. PCA, in practice, requires the probability distribution $p\left(\bm{s}\right)$ of the vertices $\bm{s}$ to calculate the covariance matrix. How does $p\left(\bm{s}\right)$ affect the projection of hypercubes? How do the resulting coordinates of the vertices change when $p\left(\bm{s}\right)$ is varied? In this section, we project hypercubes using PCA while varying the probability distribution $p\left(\bm{s}\right)$ of the vertices. Throughout the examples, we reveal the trends of the resulting projections: the leading PC corresponds to the vertices with higher probability, and PCA distorts the distance between the vertices around the origin of the projected space.
Figure 4: Orthogonal projections of four-dimensional hypercubic vertices using PCA. (a) A projection of a four-dimensional hypercube where vertices are weighted randomly. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertex. Arrows are biplot bases and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. (b) Fraction of explained variance by each PC of (a). (c) PC1 loading, and (d) PC2 loading of random weighted hypercubic vertices of (a). (e) Hamiltonian path on a four-dimensional hypercube in (a). For a different realization of random weight, see the Supplemental Material [105]. (f–j) Same as (a–e) but with bipolar distribution. Two of the vertices, $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, are more weighted than the others.
@@ IV.1 Random distribution
The biplot bases in Fig. 4(a) are drawn as arrows from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, assisting us in estimating the original binary coordinates of the vertices. Here, we abbreviate $+1$ to $+$ and $-1$ to --. For example, the projected vertex around $\begin{bmatrix}-2.0&-0.25\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}-&-&+&-\end{bmatrix}^{\top}$ and the projected vertex around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Equivalently, one can follow the hypercubic edge from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ (the origin of the biplot bases) to the vertex of interest by changing the corresponding elements indicated by the biplot bases to know the original Ising coordinate. Notice that it is not necessary to start from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ to infer the original Ising coordinate: if one knows the original coordinates of any vertex, one can infer the rest from the biplot bases, cf. Eq. (2).
@@
We then present the usage of the projection of the hypercube using PCA. In Fig. 4(e), the Hamiltonian path on the hypercube is shown following the biplot. Starting from the vertex $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, one can follow the path to the vertex $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ in the original high-dimensional space, by knowing which digit changed by following the biplot bases. These properties may be useful for understanding pathways of state transition of the target system. Notice, however, that the arrows indicating the Hamiltonian path overlap around the origin of the projected space. This overlap makes it difficult to follow the path. We address this issue in Sec. VI.
Figure 5: Orthogonal projections of four-dimensional hypercubic vertices using PCA. $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ are the most weighted vertices, $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$ are the second most weighted vertices, $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$ are the third most weighted vertices, and the rest of them are weighted randomly. (a) Projection of the hypercube by PC1 and PC2. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertices. Arrows are biplot bases and the boxes on the bottom right show the correspondence between the colors of the arrows and the original dimensions. Dashed arrows correspond to the projection of original high-dimensional vectors, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. The original coordinate is indicated around the lower right of the arrowhead as an array of filled $\blacksquare$ (indicates $+$) or empty $\square$ (indicates --) boxes. Horizontal and vertical dashed lines crossing the origin are for visual aid. (b) Same as (a) but by PC1 and PC3. (c) Fraction of explained variance by each PC. (d) PC1 loading, (e) PC2 loading, and (f) PC3 loading.

### v3

@@ Abstract
Projections of hypercubes have been applied to visualize high-dimensional binary state spaces in various scientific fields. Conventional methods for projecting hypercubes, however, face practical difficulties. Manual methods require nontrivial adjustments of the projection basis, while optimization-based algorithms limit the interpretability and reproducibility of the resulting plots. These limitations motivate us to explore theoretically analyzable projection algorithms such as principal component analysis (PCA). Here, we investigate the mathematical properties of PCA-projected hypercubes. Our numerical and analytical results show that PCA effectively captures polarized distributions within the hypercubic state space. This property enables the assessment of the asymptotic distribution of projected vertices and error bounds, which characterize the performance of PCA in the projected space. We demonstrate the application of PCA to visualize the hypercubic energy landscapes of Ising spin systems. By adding projected hypercubic edges, these visualizations reveal pathways of correlated spin flips. Our work provides a better understanding of how PCA discovers hidden patterns in high-dimensional binary data.
@@ I Introduction
Human interest in higher dimensions, particularly the fourth dimension, has been emerging since the 19th century [1, 2, 3, 4]—before the theory of relativity by A. Einstein [5, 6, 7], which is well known as the theory of four-dimensional spacetime [8]. Unlike the theory of relativity, 19th-century mathematicians, physicists, and philosophers focused on the spatial fourth dimension. Among them, C. H. Hinton has been influential because his works [9] indicate that the tesseract, or four-dimensional cube, can embody four-dimensional space. Since then high-dimensional space has been explored with hypercubes.
@@ II.2 Contribution basis
The interpretation of the contribution basis in Eq. (2) is the contribution from the $i$th dimension to the projected space. Because $b_{i}\in\left\{1,0\right\}$, the contribution basis corresponds to the projected unit vector of each dimension. Thus, by introducing the contribution basis, we can decompose the resulting projection into the contributions from each dimension as shown in Eq. (2). The remaining question is how to determine the contribution basis $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$.
Figure 2: A gallery of orthogonal projections of hypercubes. Colored arrows represent the contribution basis corresponding to the original dimensions. The boxes on the bottom right indicate the correspondence between the colors and the original dimensions. (a) An isometric projection of a cube. Notice that $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$ are overlapped. (b) An isometric projection of a tesseract. (c) A Hamming projection of a cube. (d) A Hamming projection of a tesseract. Notice that $\begin{bmatrix}1&0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1&0\end{bmatrix}^{\top}$ are overlapped. In Hamming projections (c) and (d), the contribution basis of each dimension has the same horizontal contribution. (e) A fractal projection of a six-dimensional hypercube. (f) A fractal projection of an eight-dimensional hypercube. In fractal projections (e) and (f), the contribution basis of the first half (left half) of the code are ten times longer than the rest. More projections of hypercubes are available in the Supplemental Material [109].
@@ II.3 Isometric projection
One particularly regular and symmetric projection is called isometric projection, where a Petrie polygon is projected as a regular polygon located at the periphery of the projection of a polytope [10]. Petrie polygons are equatorial polygons, lying in planes crossing the center of the object and inscribed in great circles of the circumsphere of the object. To achieve isometric projection, one needs the projected unit vector of the $i$th dimension to form consecutive edges of the projected Petrie polygon, i.e., $\bm{e}_{i}$ form a Petrie polygon together with a reversed vector $-\bm{e}_{i}$. To project the Petrie polygon as a regular polygon, the contribution basis are determined as $\bm{e}_{i}=\begin{bmatrix}\cos\left(\theta_{i}\right)&\sin\left(\theta_{i}\right)\end{bmatrix}^{\top}$, where $\theta_{i}=\frac{\uppi}{N}\left(i-1\right)+\phi$, with a constant $\phi$. The resulting isometric projections and vectors $\bm{e}_{i}$ used to create them are shown in Figs. 2(a) and 2(b), where Petrie polygons are projected as a regular hexagon [Fig. 2(a)] and octagon [Fig. 2(b)]. We determine the angle $\phi=\frac{\uppi}{2N}$ for symmetry of the projections. In isometric projections, all edges of the hypercube are drawn with the same length.
@@ II.4 Hamming and fractal projections
By modifying the contribution basis $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ for projection, one can view a high-dimensional hypercube from various angles. Several methods reflecting the cross-section of the hypercube have already been suggested [10]. Here, we introduce two other orthogonal projections of hypercubes used in sciences.
@@ II.5 Exchanging the labels of vertices
Here, we show a visual understanding of this conversion through contribution basis. By reversing the direction of the contribution basis, one can swap the projected coordinates of the rest, cf. Figs. 2(a) and 2(b), or Figs. 3(a) and 3(b). The dimension of the reversed contribution basis corresponds to the dimension where there is a $1$ in $\bm{b}^{\prime}\oplus\bm{b}^{\prime\prime}$. Visualizing the projected vector provides a way to understand exchanging the labels via reversing the direction of unit vectors.
Figure 3: Exchanging the labels of vertices by reversing the direction of the contribution basis of corresponding digits. The boxes on the bottom right indicate the correspondence between the colors and the dimensions. (a) By reversing the contribution basis for the second digit, one can exchange the cubic labels of pairs of vertices, $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1\end{bmatrix}^{\top}$, and $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Compare this with Fig. 2(a). (b) Obtaining different labels for a tesseract by swapping two contribution basis. The first and fourth unit vectors are reversed, cf. Fig. 2(b).
@@ II.6 The limitation of manual orthogonal projections
Although one can create reproducible and interpretable (and even visually label-exchangeable) two-dimensional projections of hypercubes in the abovementioned way, the disadvantage is that one must manually determine the contribution basis of each dimension. For low-dimensional hypercubes, it is manageable to adjust the contribution basis oneself to make the projections easy to interpret. Still, it is nontrivial to decide on contribution basis for high-dimensional hypercubes. Thus, automatic and unsupervised projection methods are demanded, especially for projections of high-dimensional hypercubes. How can one project a hypercube with such a strategy?
@@ III.2 PCA and biplot
From now on, we call $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ biplot basis. Notice that, when the data points are binary vertices $\bm{b}\in\left\{1,0\right\}^{N}$ of a system, the biplot basis exactly match the basis of the orthogonal projection of a hypercube, while when the data points are Ising vertices $\bm{s}\in\left\{+1,-1\right\}^{N}$, the biplot basis with doubled magnitudes exactly match the contribution basis. This difference arises from the difference in the length of hypercubic edges: they are one for binary vertices and two for Ising vertices. This is why we add factor $2$ in Eq. (10). By performing PCA and plotting biplot basis, one can obtain a reproducible, interpretable, and automatic projection of a hypercube.
@@ IV Hypercubic PCA
In the previous Sec. III, we introduced PCA as a method to project hypercubes and presented the interpretation of the projection using biplot basis. PCA, in practice, requires the probability distribution $p\left(\bm{s}\right)$ of the vertices $\bm{s}$ to calculate the covariance matrix. How does $p\left(\bm{s}\right)$ affect the projection of hypercubes? How do the resulting coordinates of the vertices change when $p\left(\bm{s}\right)$ is varied? In this section, we project hypercubes using PCA while varying the probability distribution $p\left(\bm{s}\right)$ of the vertices. Throughout the examples, we reveal the trends of the resulting projections: the leading PC corresponds to the vertices with higher probability, and PCA distorts the distance between the vertices around the origin of the projected space.
Figure 4: Orthogonal projections of four-dimensional hypercubic vertices using PCA. (a) A projection of a four-dimensional hypercube where vertices are weighted randomly. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertex. Arrows are biplot basis and the boxes on the bottom right indicate the correspondence between the colors of the arrows and the original dimensions. (b) Fraction of explained variance by each PC of (a). (c) PC1 loading, and (d) PC2 loading of random weighted hypercubic vertices of (a). (e) Hamiltonian path on a four-dimensional hypercube in (a). For a different realization of random weight, see the Supplemental Material [109]. (f–j) Same as (a–e) but with bipolar distribution. Two of the vertices, $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$, are more weighted than the others.
@@ IV.1 Random distribution
The biplot basis in Fig. 4(a) are drawn as arrows from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, assisting us in estimating the original binary coordinates of the vertices. Here, we abbreviate $+1$ to $+$ and $-1$ to --. For example, the projected vertex around $\begin{bmatrix}-2.0&-0.25\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}-&-&+&-\end{bmatrix}^{\top}$ and the projected vertex around $\begin{bmatrix}-0.4&1.7\end{bmatrix}^{\top}$ corresponds to the vertex $\begin{bmatrix}+&-&+&+\end{bmatrix}^{\top}$. Equivalently, one can follow the hypercubic edge from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ (the origin of the biplot basis) to the vertex of interest by changing the corresponding elements indicated by the biplot basis to know the original Ising coordinate. Notice that it is not necessary to start from $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ to infer the original Ising coordinate: if one knows the original coordinates of any vertex, one can infer the rest from the biplot basis, cf. Eq. (2).
@@
We then present the usage of the projection of the hypercube using PCA. In Fig. 4(e), the Hamiltonian path on the hypercube is shown following the biplot. Starting from the vertex $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$, one can follow the path to the vertex $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ in the original high-dimensional space, by knowing which digit changed by following the biplot basis. These properties may be useful for understanding pathways of state transition of the target system. Notice, however, that the arrows indicating the Hamiltonian path overlap around the origin of the projected space. This overlap makes it difficult to follow the path. We address this issue in Sec. VI.
@@ IV.3 Sexapolar distribution
Figure 5: Orthogonal projections of four-dimensional hypercubic vertices using PCA. $\begin{bmatrix}-&-&-&-\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$ are the most weighted vertices, $\begin{bmatrix}-&-&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&-&-\end{bmatrix}^{\top}$ are the second most weighted vertices, $\begin{bmatrix}-&+&-&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&-&+&-\end{bmatrix}^{\top}$ are the third most weighted vertices, and the rest of them are weighted randomly. (a) Projection of the hypercube by PC1 and PC2. Red filled circles are the vertices and lines are the edges of the hypercube. The magnitude of weight is proportional to the area of the vertices. Arrows are biplot basis and the boxes on the bottom right show the correspondence between the colors of the arrows and the original dimensions. Dashed arrows correspond to the projection of original high-dimensional vectors, $\begin{bmatrix}-&+&+&+\end{bmatrix}^{\top}$ and $\begin{bmatrix}+&+&+&+\end{bmatrix}^{\top}$. The original coordinate is indicated around the lower right of the arrowhead as an array of filled $\blacksquare$ (indicates $+$) or empty $\square$ (indicates --) boxes. Horizontal and vertical dashed lines crossing the origin are for visual aid. (b) Same as (a) but by PC1 and PC3. (c) Fraction of explained variance by each PC. (d) PC1 loading, (e) PC2 loading, and (f) PC3 loading.